# IJCAI–25 Formatting Instructions

### v1

@@ 5.2 NMT Architectures
We consider two RANMT architectures. The first is an in-house implementation of NFA (Bulte and Tezcan, 2019), which extends a basic encoder-decoder MT architecture by prepending the target side of TM examples to the source text on the encoder side. As is standard, the retrieved examples used in training are selected based on their Levenshtein distance to the source text; at most one similar example is used. This model thus also has the ability to perform inference from scratch, without using any retrieved example.
@@ 6 Results
We first report the results of controlled experiments, in order to contrast mono-lingual vs. cross-lingual retrieval strategies. We then analyze the results of a larger-scale setting, where we can only exploit monolingual data.
@@
$\operatorname{Lev}$uparrow xsim $\downarrow$ $\operatorname{Lev}$uparrow xsim $\downarrow$
@@
Table 1: Average retrieval scores (xsim error rate and average Levenshtein similarity), averaged for all domains.
@@ 6.1 Retrieval Scores
Table 1 stores the retrieval scores (§4.2) for the small-scale experiments. For en-fr, source-side fuzzy matching yields more similar segments than CLIR techniques, a gap that vanishes after fine-tuning with MAE for dense+bow and even more so for LaBSE. For de-en, the source-side fuzzy matching baseline is much worse, revealing a large gap with the oracle condition, which might be due to residual misalignments or to morphological differences between German and English. [^7] Inflection phenomena in German create multiple variants for each lexeme, making exact matches at the word level much less likely than in English. All CLIR techniques widely outperform the monolingual settings (fuzzy-src and fuzzy-bt). Here, however we do not see much gain from fine-tuning, perhaps owing to a smaller fune-tuning dataset. As expected, fine-tuning with a lexical loss is slightly detrimental to the sentence-level retrieval scores (xsim error) for both dense+bow and LaBSE. Training dense+bow along with a bag-of-word loss (eq. (10)) effectively improves the retrieval scores.
@@ 6.2 Translation Scores
We use both NLLB and NFA without any retrieved examples as our machine translation baselines. Their corresponding BLEU scores, averaged across domains, are respectively 39.0 (en-fr) and 33.8 (de-en) for NLLB and 48.0 (en-fr) and 38.2 (de-en) for NFA. Details per domain are in Table 4.
TM3-LevT NFA
test en-fr test de-en test en-fr test de-en
BLEU COMET BLEU COMET BLEU COMET BLEU COMET
sans exemple - - - - 48.0 87.1 38.2 79.8
fuzzy-gold (oracle) 45,3 51,7 31,3 -8,6 52,2 87,8 44,7 82,2
fuzzy-src 43,8 48,4 28,8 -10,3 51,3 87,5 41,9 81,7
fuzzy-bt 40,5 43,8 22,0 -22,9 49,9 87,2 38,9 81,3
dense 41,5 42,2 - - 49,9 87,1 - -
dense+bow 42,4 44,1 28,3 -11,6 50,5 87,3 40,0 80,9
ft-dense+bow-MSE 43,1 46,0 29,0 -12,7 51,0 87,5 41,6 81,5
ft-dense+bow-MAE 43,2 46,3 29,1 -12,6 51,0 87,4 41,8 81,4
ft-dense+bow-Rank 42,1 44,3 26,0 -17,9 50,7 87,3 41,8 80,9
LASER 42,5 44,3 30,4 -7,7 49,9 87,3 41,1 81,3
LaBSE 43,1 45,5 29,8 -9,5 50,5 87,3 41,5 81,4
ft-LaBSE-MSE 43,7 47,8 27,7 -12,5 51,2 87,5 41,6 81,6
ft-LaBSE-MAE 43,6 47,7 29,1 -9,2 51,2 87,5 42,0 81,5
ft-LaBSE-Rank 43,3 47,1 28,4 -12,2 51,1 87,5 41,8 81,6
Table 2: Average translation scores: TM3-LevT and NFA, all domains.
@@
Table 4: Per domain BLEU scores for NFA systems and NLLB. The contamination rate (top line) for NFA is the percentage of overlap of its training data with the test set. Contaminated domains for NLLB (used in back-translation) are underlined.
@@
Translation results achieved with TM3-LevT are in Table 2. The difference between the oracle (fuzzy-gold) and the baseline (fuzzy-src) gives an idea of the quality of the monolingual matches: the quality is pretty high in en-fr, owing to the cleaning procedure used for these data, more noisy in de-en, with a gap of about $2.5$ BLEU points. In comparison, using artificial back-translated source queries performs much worse than using natural texts. This difference is much larger than what was expected from the retrieval scores of Table 1.
@@
The corresponding results for NFA are in Table 2. This model is a much stronger baseline than TM3-LevT and the margin of progress (up to fuzzy-gold) is dimmer. Yet, we see that (a) retrieving examples directly on the target side (with dense, LASER, and LaBSE) already improves the example-free baseline by about $3$ BLEU points; (b) an extra step of fine-tuning of cross-lingual retrievers (with MSE and MAE) enables us to fully close the gap with source-side matching for both metrics and language pairs. In comparison, back-translating the monolingual target data also improves our baseline systems, but by a much smaller margin.
@@
Per-domain BLEU scores are in Tables 3 (TM3-LevT) and 4 (NFA). The domains for which we expect CLIR methods to work well need to satisfy two conditions: (a) a large enough set of examples to retrieve from, so that the choice the retriever actually makes a difference; (b) a small diversity of texts in that domain, so that retrieved instances are sufficiently similar to the input. This is well reflected in Table 3 (TM3-LevT) where CLIR techniques (with fine-tuning in en-fr, without in de-en) match the performance of the baseline in all domains, and even outpeform it in specialized domains such as ECB (finance), JRC (law), KDE (information technology) and Wikipedia. We do not see the same gains in Ubuntu (very small TM) or Europarl (large TM, with a diverse set of texts).
@@
The standard contrastive loss of equation (9), when applied in MT, primarily seeks to identify parallel sentences within collections of monolingual segments. Yet, RANMT requires examples that are lexically similar to the source sentences. The bag-of-word loss of equation (10) effectively increases the rate of lexically similar sentences, thus enhancing the translation scores. As for our three fine-tuning strategies, MSE and MAE (eq. (5)) yield comparable performances. The per-domain analysis further highlights their consistent behavior. As for Rank (eq. (8)), we observe that the rank-based loss consistently underperforms the other two, with a few exceptions (PHP for NFA; PHP, Wiki and KDE (de-en) for TM3-LevT).
@@
We identified two sources of data contamination. First, NFA training data partly overlaps with some of our test sets, in varying proportion (from 0 to 100) across domains (see Table 4). The intuition is that such contamination should make this model less sensitive to the choice of good examples. In the per-domain results of table 4, this sensitivity is assessed by the gap between example-free and fuzzy-gold scores. As this gap is large for most domains (with the exception of Europarl and News in en-fr, and Sub in de-en), we can still clearly assess the benefits of cross-lingual retrieval for the large majority of domains. Second, the training data of NLLB, used for back-translation, also partly overlaps with our test, which may boost the efficiency of the fuzzy-bt baseline. Even for these domains (e.g., ECB, EMEA, JRC-Acquis), cross-lingual methods still achieve much higher BLEU scores than back-translation.
Figure 1: Evolution of BLEU and COMET scores of TM3-LevT (top) and NFA (bottom) when growing the retrieval pool.
6.3 A Large-scale Experiment
@@
We use the models trained on the en-fr dataset to retrieve segments from the large monolingual Wikipedia dataset (see §4.1), using the Wikipedia test already used in the controled experiments set to assess performance. We use the same filtering threshold as for the controlled experiments to keep the settings comparable and ensure that the only difference is the increased size of the retriveal set. We only consider the variants of LaBSE, which achieve the best results in the small data condition.
Wikipedia is a challenging domain, due to the variety of topics that are addressed in the encyclopedia and the high level of formality of the text. NLLB already represents a strong baseline for that domain. Results are in Table 5. The first observation is that increasing the retrieval set (from 200K segments in the controlled experiments – Tables 3 and 4 – to 45M segments) yields significant performance boosts across the board (e.g, +1.6 BLEU for NFA and fuzzy-bt, +0.4 BLEU points for NFA and LaBSE). With an additional lexical fine-tuning step, the benefits of the cross-lingual approach are again very clear, with large BLEU differences relative to the small data condition:+3.9 for ft-LaBSE-MAE with TM3-LevT, +3.4 when using NFA.
TM3-LevT NFA
BLEU COMET BLEU COMET
no example - - 36.4 84.8
fuzzy-bt 35.6 76.7 38.6 84.8
LaBSE 37.7 77.2 40.1 85.0
ft-LaBSE-MSE 38.3 77.2 40.9 84.8
ft-LaBSE-MAE 38.7 77.6 41.2 84.9
ft-LaBSE-Rank 38.5 77.5 40.8 85.0
Table 5: BLEU and COMET scores with systems TM3-LevT and NFA with the Wikipedia retrieval pool.
@@
To better analyse the effect of the retrieval pool size, we progressively increase the Wikipedia retrieval set by batches of $500$K sentences from $500$K to $45$M, and compute the corresponding translation scores for each of the retrieval systems. Figure 1 reports the evolution of the translation metrics for LaBSE, its fine-tuned variants and fuzzy-bt. Overall, for all models, the benefit of growing the retrieval pool yields an almost linear BLEU increase. COMET variations are much smaller, especially for NFA where it remains nearly constant for our best setups. For both metrics, ft-LaBSE-MAE outperforms the other setups for all data sizes. fuzzy-bt consistently lags behind CLIR alternatives, especially with TM3-LevT.
We also monitor in table 6 two additional scores: the retrieval rate (percentage of segments for which at least one example is found) and the average $\operatorname{Lev}$ similarity between the reference and its closest retrieved neighbor. For all retrieval methods, we see large increases, meaning a larger number of retrieved examples is retrieved, and that their average quality (i.e their similarity to the reference) also improves, highliting the usefulness of monolingual data for this domain.
@@
RR $\uparrow$ $\operatorname{Lev}$uparrow RR $\uparrow$ $\operatorname{Lev}$uparrow
@@
Table 6: Comparison of the retrieval rates (RR) and average Levenshtein similarity ($\operatorname{Lev}$) on the Wikipedia test set of the systems retrieving in the TM vs. in the large monolingual corpus (mono).

### v2

@@ 5.2 NMT Architectures
We consider three RANMT architectures. The first is an in-house implementation of NFA (Bulte and Tezcan, 2019), which extends a basic encoder-decoder MT architecture by prepending the target side of TM examples to the source text on the encoder side. As is standard, the retrieved examples used in training are selected based on their Levenshtein distance to the source text; at most one similar example is used. This model thus also has the ability to perform inference from scratch, without using any retrieved example.
@@
The third architecture is EuroLLM-9B (Martins et al., 2025), specifically designed and trained for multilingual tasks such as Machine Translation. We fine-tune the EuroLLM-9B base model on the en-fr and de-en training corpora using a prompt that only contains relevant examples in the target language. A discussion of the adaption of EuroLLM for this setup is in Appendix D.
@@ 6 Results
We first run controlled experiments serving two main purposes: (a) to compare the quality of various retrieval models (§ 6.1); (b) to evaluate the gap between the baseline source-side retrieval and our proposed CLIR-based approaches (§ 6.2), an analysis that crucially requires parallel datasets. These experiments use two language pairs (en-fr and de-en). This configuration is somewhat ideal, as CLIR methods are primarily meant for cases where only monolingual data are available, as illustrated by the experiments in § 6.3.
@@
$\operatorname{Lev}$ $\uparrow$ xsim $\downarrow$ $\operatorname{Lev}$ $\uparrow$ xsim $\downarrow$
@@
Table 1: Average retrieval scores (xsim error and Levenshtein similarity), averaged accross domains.
@@ 6.1 Retrieval Scores
Table 1 stores the retrieval scores (§4.2) for the small-scale experiments. For en-fr, source-side fuzzy matching yields more similar segments than CLIR techniques, a gap that vanishes after fine-tuning with MAE for dense+bow and even more so for LaBSE. For de-en, the gap between the source-side fuzzy matching baseline and the oracle condition is much larger than for en-fr, which might be due to residual misalignments or to morphological differences between German and English. [^10] Inflectional processes in German create multiple variants for each lexeme, making exact matches at the word level much less likely in German than in English. All CLIR techniques widely outperform the monolingual settings (fuzzy-src and fuzzy-bt). For this language pair, however we do not see much gain from fine-tuning, perhaps owing to a smaller fune-tuning dataset. As expected, fine-tuning with a lexical loss is always slightly detrimental to the sentence-level retrieval scores (xsim error) for both dense+bow and LaBSE. Training dense+bow along with a bag-of-word loss (eq. (11)) effectively improves the retrieval scores.
@@ 6.2 Translation Scores
We use NLLB, NFA and EuroLLM without any retrieved examples as machine translation baselines. Their corresponding BLEU scores, averaged across domains, are respectively 39.0 (en-fr) and 33.8 (de-en) for NLLB and 48.0 (en-fr) and 38.2 (de-en) for NFA. Details per domain are in Table 4.
TM3-LevT NFA EuroLLM
test en-fr test de-en test en-fr test de-en test en-fr test de-en
BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET
no example - - - - 48.0 87.1 38.2 79.8 44.4 86.3 40.7 82.6
fuzzy-gold (oracle) 45.3 51.7 31.3 -8.6 52.2 87.8 44.7 82.2 52.6 88.1 46.8 83.9
fuzzy-src 43.8 48.4 28.8 -10.3 51.3 87.5 41.9 81.7 51.6 87.8 41.6 83.0
fuzzy-bt 40.5 43.8 22.0 -22.9 49.9 87.2 38.9 81.3 46.4 86.9 41.8 82.9
dense 41.5 42.2 - - 49.9 87.1 - - 50.4 87.4 - -
dense+bow 42.4 44.1 28.3 -11.6 50.5 87.3 40.0 80.9 51.1 87.5 44.1 83.0
ft-dense+bow-MSE 43.1 46.0 29.0 -12.7 51.0 87.5 41.6 81.5 51.3 87.8 45.1 83.3
ft-dense+bow-MAE 43.2 46.3 29.1 -12.6 51.0 87.4 41.8 81.4 51.5 87.7 45.2 83.4
ft-dense+bow-Rank 42.1 44.3 26.0 -17.9 50.7 87.3 41.8 80.9 50.4 87.5 43.6 83.1
LASER 42.5 44.3 30.4 -7.7 49.9 87.3 41.1 81.3 50.9 87.4 44.9 83.1
LaBSE 43.1 45.5 29.8 -9.5 50.5 87.3 41.5 81.4 51.5 87.6 45.1 83.2
ft-LaBSE-MSE 43.7 47.8 27.7 -12.5 51.2 87.5 41.6 81.6 51.6 87.8 44.2 83.2
ft-LaBSE-MAE 43.6 47.7 29.1 -9.2 51.2 87.5 42.0 81.5 51.5 87.8 44.9 83.3
ft-LaBSE-Rank 43.3 47.1 28.4 -12.2 51.1 87.5 41.8 81.6 51.3 87.7 44.8 83.2
Table 2: Average translation scores: TM3-LevT, NFA, EuroLLM, all domains. Significant results (p=0.05) w.r.t. LaBSE are marked with . Best CLIR results are in bold. When fuzzy-src is not outperformed, its score is in bold.
@@
Table 4: Per domain BLEU scores for NLLB and NFA systems. Contaminated domains for NLLB (used in back-translation) are underlined. The contamination rate (top line) for NFA is the percentage of overlap of its training data with the test set.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
no example 49.9 47.7 39.2 45.7 55.5 40.9 34.0 48.1 42.0 41.3 44.2 41.6 23.3 54.4 52.9 31.1
fuzzy-gold 63.7 68.4 39.5 65.2 66.2 51.1 33.8 52.1 41.9 48.9 48.2 50.4 25.6 63.6 62.2 32.2
fuzzy-src 62.6 66.9 39.5 63.3 64.9 49.0 33.7 51.6 41.7 47.4 47.0 43.3 23.6 55.1 54.3 31.8
fuzzy-bt 52.0 54.1 39.3 49.1 56.5 43.2 33.8 50.4 41.5 45.0 46.1 42.5 23.6 54.8 56.4 31.9
LASER 63.2 65.3 39.4 62.1 65.5 46.1 33.8 51.0 41.7 45.8 46.4 46.6 24.4 62.4 59.9 31.4
LaBSE 63.5 66.1 39.4 63.3 65.8 48.4 33.8 51.2 41.6 46.4 46.5 47.2 23.7 62.7 60.3 31.4
ft-LaBSE-MSE 62.8 66.3 39.4 63.2 64.9 50.0 33.8 51.7 41.7 46.8 47.3 46.6 24.1 60.0 58.7 31.5
ft-LaBSE-MAE 62.8 65.9 39.4 63.3 65.1 49.6 33.8 51.5 41.7 46.8 47.1 47.6 24.2 61.3 59.9 31.6
ft-LaBSE-Rank 62.4 65.3 39.4 62.7 64.6 49.3 33.8 51.5 41.7 47.3 46.7 47.5 23.8 61.1 60.2 31.5
Table 5: Per domain BLEU scores for EuroLLM.
@@
Translation results for TM3-LevT are in Table 2 (left part). The difference between the oracle (fuzzy-gold) and the baseline (fuzzy-src) gives an idea of the quality of the monolingual matches: the quality is pretty high in en-fr, owing to the cleaning procedure used for these data, more noisy in de-en, with a gap of about $2.5$ BLEU points. In comparison, using artificial back-translated source queries performs much worse than using natural texts. This difference is much larger than what was expected from the retrieval scores of Table 1.
@@
The corresponding results for NFA are in Table 2 (middle part). This model is a much stronger baseline than TM3-LevT and the progress margins (with respect to fuzzy-gold) is dimmer. Yet, we see that (a) retrieving examples directly on the target side (with dense, LASER, and LaBSE) improves the example-free baseline by about $3$ BLEU points; (b) further fine-tuning the cross-lingual retrievers (with MSE and MAE) fully closes the gap with source-side matching for both metrics and language pairs. We observe that back-translating the monolingual target data also improves baseline systems, albeit by a much smaller margin.
EuroLLM
Results for the fine-tuned version of EuroLLM are in the right part of Table 2. This system turns out to be the most effective system overall, slightly outperforming NFA with $k=3$ TM examples. A first observation is that this model is particularly strong for translating into English (from German) where using a CLIR outperforms all baselines. For this model, the averaged gains achieved by fine-tuning the cross-lingual retrievers are very small, and vary accross domains (details in Table 5). For this system, the gap between the LaBSE system and the oracle condition (fuzzy-gold) is already small in the two language pairs, showcasing the very strong built-in cross-lingual abilities of this model; yet, the difference that remains suggests that lexical similarity should still matter.
@@
Per-domain BLEU scores are in Tables 3 (TM3-LevT), 4 (NFA), and 5 (EuroLLM). The domains for which we expect CLIR methods to work well need to satisfy two conditions: (a) a large enough set of examples to retrieve from, so that the choice the retriever actually makes a difference; (b) a small diversity of texts in that domain, so that retrieved instances are sufficiently similar to the input. This is well reflected for instance in Table 3 (TM3-LevT) where CLIR techniques (with fine-tuning in en-fr, without in de-en) match the performance of the baseline in all domains, and even outpeform it in specialized domains such as ECB (finance), JRC (law), KDE (information technology) and Wikipedia. We do not see the same gains in Ubuntu (very small TM) or Europarl (large TM, with a diverse set of texts).
@@
Regarding EuroLLM, we note that performing CLIR with the default version of LaBSE already yields near optimal performance for many domains. Fine-tuning has here a reduced effect on BLEU scores, sometimes slightly improving (e.g., for en-fr: EMEA, KDE, PHP, Wiki, UBU) or, more rarely, decreasing (e.g., for en-fr, ECB, JRC) performance. These variations highlight the fact that optimal translation scores ultimately also depend on the retrieval pool and the associated example quality.
@@
The standard contrastive loss of eq. (10), when applied in MT, primarily seeks to identify parallel sentences within collections of monolingual segments. Yet, RANMT requires examples that are lexically similar to the source sentences. The bag-of-word loss of eq. (11) effectively increases the rate of lexically similar sentences, thus enhancing the translation scores. As for our three fine-tuning strategies, MSE and MAE (eq. (5)) yield comparable performances. The per-domain analysis further highlights their consistent behavior. As for Rank (eq. (9)), we observe that the rank-based loss consistently underperforms the other two, with a few exceptions (PHP for NFA; PHP, Wiki and KDE (de-en) for TM3-LevT).
@@
We identify two sources of data contamination. First, NFA training data partly overlaps with some of our test sets, in varying proportion (from 0 to 100) across domains (see Table 4). The intuition is that such contamination should make this model less sensitive to the choice of good examples. In the per-domain results of Table 4, this sensitivity is assessed by the gap between example-free and fuzzy-gold scores. As this gap is large for most domains (except Europarl and News in en-fr, and Sub in de-en), we can still clearly assess the benefits of cross-lingual retrieval for the large majority of domains. Second, the training data of NLLB, used for back-translation, also partly overlaps with our test, which may boost the efficiency of the fuzzy-bt baseline. Even for these domains (e.g., ECB, EMEA, JRC-Acquis), cross-lingual methods still achieve much higher BLEU scores than back-translation.
Figure 1: BLEU and COMET scores with a growing retrieval pool: TM3-LevT (top), NFA (middle) and EuroLLM (bottom).
6.3 Large-scale Experiments
@@
We use the models trained on the en-fr [^11] A comparable experiment, targeting translation from English into Ukrainian is in Appendix H; these results also highlight the benefits of CLIR over monolingual retrieval in parallel data, especially when using a fine-tuned version of LaBSE. dataset to retrieve segments from the large monolingual Wikipedia dataset (see §4.1), using the Wikipedia test already used in the controlled experiments set to assess performance. We use the same filtering threshold as for the controlled experiments to keep the settings comparable and ensure that the only difference is the increased size of the retrieval set. We only consider the variants of LaBSE, which achieve the best results in the small data condition.
Wikipedia is a challenging domain, due to the variety of topics that are addressed in the encyclopedia and the high level of formality of the text. NLLB already represents a strong baseline for that domain. Results are in Table 6. The first observation is that increasing the retrieval set (from 200K segments in the controlled experiments – Tables 3 and 4 – to 45M segments) yields significant performance boosts across the board (e.g, +1.6 BLEU for NFA and fuzzy-bt, +0.4 BLEU points for NFA and LaBSE). With an additional lexical fine-tuning step, the benefits of the cross-lingual approach are again very clear, with large BLEU differences relative to the small data condition:+3.9 for ft-LaBSE-MAE with TM3-LevT, +3.4 when using NFA, +2 using EuroLLM.
TM3-LevT NFA EuroLLM
BLEU COMET BLEU COMET BLEU COMET
no example - - 36.4 84.8 44.2 86.3
fuzzy-bt 35.6 76.7 38.6 84.8 47.1 86.4
LaBSE 37.7 77.2 40.1 85.0 48.8 86.5
ft-LaBSE-MSE 38.3 77.2 40.9 84.8 49.0 86.3
ft-LaBSE-MAE 38.7 77.6 41.2 84.9 49.2 86.3
ft-LaBSE-Rank 38.5 77.5 40.8 85.0 48.6 86.3
Table 6: BLEU and COMET scores with systems TM3-LevT, NFA and EuroLLM with the Wikipedia retrieval pool.
@@
To better analyse the effect of the retrieval pool size, we progressively increase the Wikipedia retrieval set by batches of 500500K sentences from 500500K to 4545M, and compute the corresponding translation scores for each retriever. Figure 1 displays the evolution of the translation metrics for LaBSE, its fine-tuned variants and fuzzy-bt. Overall, for all models, the benefit of growing the retrieval pool yields an almost linear BLEU increase. COMET variations are much smaller, especially for NFA and EuroLLM where it remains nearly constant. For both metrics, ft-LaBSE-MAE outperforms the other setups for all data sizes in most cases. fuzzy-bt consistently lags behind CLIR alternatives, especially with TM3-LevT.
We also monitor in Table 7 two additional scores: the retrieval rate (percentage of segments for which at least one example is found) and the average $\operatorname{Lev}$ similarity between the reference and its closest retrieved neighbor. For all retrieval methods, we see large increases, meaning a larger number of examples is retrieved and that their average quality (i.e, their similarity to the reference) also improves, highliting the usefulness of monolingual data for this domain.
@@
$R$ $\uparrow$ $\operatorname{Lev}$ $\uparrow$ $R$ $\uparrow$ $\operatorname{Lev}$ $\uparrow$
@@
Table 7: Comparison of the retrieval rates ($R$) and average Levenshtein similarity ($\operatorname{Lev}$) on the Wikipedia test set of the systems retrieving in the TM vs. in the large monolingual corpus (mono).
The translation examples in Appendix F illustrate the benefits of enlarging the retrieval pools, notably because they provide longer lexical matches, thereby improving the generated text.